### v1

# Long distance propagation of wave beams in paraxial regime

Abstract
This paper concerns the propagation of high frequency wave-beams in highly turbulent atmospheres. Using a paraxial model of wave propagation, we show in the long-distance weak-coupling regime that the wavefields are approximately described by a complex Gaussian field whose scintillation index is unity. This provides a model of the speckle formation observed in many practical settings. The main step of the derivation consists in showing that closed-form moment equations in the Itô-Schrödinger regime are still approximately satisfied in the paraxial regime. The rest of the proof is then an extension of results derived in [5].
Keywords: Wave propagation in random media; Paraxial regime; Itô-Schrödinger regime; Scintillation; Gaussian conjecture
1 Introduction
A widely accepted conjecture in the physical literature states that high frequency waves eventually become complex circular Gaussian distributed after propagating over long distances through randomly varying media [21, 26]. The real and imaginary parts of the wave field then become independent Gaussian distributed. The wave intensity follows an exponential distribution asymptotically, implying a scintillation index converging to unity. This forms a model for speckle patterns consistent with experimental observations [1, 8]. This conjecture is not supported mathematically for general models of wave propagation based on Helmholtz equations. The highly complex nature of wave propagation in random media makes mathematical analyses of asymptotic models inherently difficult; see [3, 13] for a review of existing results.
A simpler setting is that of the paraxial approximation, where the wave-field retains its beam-like structure for long distances while backscattering is neglected; see [2, 15] for theoretical justifications of such a regime. A further approximation is the Itô-Schrödinger white noise limit of the paraxial model, which using Itô calculus, is characterized by closed form PDEs for statistical moments of all orders; see [10] for an analysis of the white noise model and [11] for a derivation starting from the paraxial approximation.
While analytical solutions to these PDEs are not known in general, they form an important tool to carry out long-distance asymptotic analyses. The structure of these PDEs up to the fourth moment was used in a series of articles [16, 17, 18, 14, 19] to understand scintillation of the energy of the wavefield.
The Gaussian conjecture was proved for the Itô-Schrödinger model of wave propagation in the scintillation (weak-coupling) regime in [5]. The derivation is based on a careful analysis of these PDEs for statistical moments of arbitrary order. That work identifies two scaling limits. In the first one called the kinetic regime, random fluctuations with small amplitude accumulate over medium-long distances of propagation so that the effective turbulence strength is of order $O(1)$. In the longer-distance, second scaling called the diffusive regime, the cumulative strength of turbulence is larger than $O(1)$. In the latter scaling, the mean wave-field vanishes while its higher statistical moments are indeed consistent with those of a circular Gaussian random variable. The observed wavefield intensity then follows an exponential distribution, characterising the speckle patterns observed in experiments.
From a technical standpoint, the analysis of such moments in [5] is primarily carried out in the Fourier domain, with the error controlled in the total variation sense. The main advantage of this topology is that it allows us to obtain error estimates in the physical domain in the uniform sense. Other related scaling limits have also been obtained for a Fourier-domain representation of a phase-compensated wavefield in [4, 22].
We note that the Gaussian conjecture is not always true, even in the Itô-Schrödinger regime. It has been shown that in the so-called spot dancing regime, the wave field approximately follows a Rice-Nakagami distribution instead for Gaussian incident beams [12, 10]. Such a regime is characterized by relatively narrow beams travelling through very strong turbulence over relatively short distances. The wave field then experiences significant beam wander with the overall profile of the beam diffracting as in the uniform medium. In contrast, the scintillation regime is characterized by broad incident beams propagating through weak turbulence over long distances. In such a regime, the wavefield decorrelates rapidly along the lateral cross-section of the beam, which is consistent with speckle patterns observed in physical experiments.
The main objective of this paper is to prove the Gaussian conjecture in the scintillation regime starting from the paraxial approximation of the scalar Helmholtz equation. As in the derivation for the Itô-Schrödinger equation in [5], this requires sufficiently rapid decorrelations of the random fluctuations along the direction of propagation in the rescaled variables. The presence of long range correlations leads to considerably different regimes [20, 7]. These requirements are encoded in the following scaling of wave propagation.
The starting microscopic model of wave-beam propagation in turbulent media is the following scalar Helmholtz equation:
$\Big{(}\partial^{2}_{z}+\Delta_{x}+k_{0}^{2}n^{2}(z,x)\Big{)}p(z,x)=-u_{0}(x)\delta_{0}(z)\,,$ (1)
where $k_{0}$ is the carrier wave number of the source, $z\in\mathbb{R}$ is depth and $x\in\mathbb{R}^{d}$ for $d\geq 1$ denotes transverse spatial variables. Suppose $n^{2}(z,x)=1+\nu(z,x)$ where $\nu$ is a random coefficient modeled as a mean zero Gaussian random process. We consider a high frequency long propagation scaling given by
$z\to\frac{\eta z}{\varepsilon\theta},\qquad x\to x,\qquad k_{0}\to\frac{k_{0}}{\theta},\qquad\nu\to\frac{\varepsilon^{\frac{1}{2}}\theta^{\frac{3}{2}}}{\eta^{\frac{3}{2}}}\nu\qquad u_{0}\to u_{0}^{\theta},$ (2)
with $0<\theta\ll 1$. As in [5], we consider the weak coupling regime $0<\varepsilon\ll 1$ in two sub-regimes, the kinetic regime when $\eta=1$ and the diffusive regime when $\eta\ll 1$. That the latter regime corresponds to long-distance propagation (optically thick media) may be obtained from the fact that distance of propagation times random strength (proportional to $\nu^{2}$) is $\eta/\varepsilon\times\varepsilon/\eta^{3}=\eta^{-2}\gg 1$. We will see below that $\varepsilon$ models the width of the incident beam $u_{0}$. For technical reasons as in [5], we need a large scale separation between $\varepsilon$ and $\eta$ in the diffusive regime and assume that $\eta=(\log|\log\varepsilon|)^{-1}$.
The results obtained in this paper hold in the kinetic regime so long as both $\varepsilon\to 0$ and $\theta\to 0$. In the diffusive regime, we will need in addition $\theta e^{\frac{C}{\eta^{2}}}/\eta^{2}$ to be controlled and asymptotically small. In order to simplify notation and not carry through the small parameters $(\theta,\varepsilon,\eta)$, we assume that $\varepsilon=\varepsilon(\theta)$ as a function (continuous from $(0,\frac{1}{2}]$ to $(0,\frac{1}{2}]$) such that $\varepsilon(\theta)\to 0$ as $\theta\to 0$. Therefore, $\eta=\eta(\theta)$ implicitly. We also assume for concreteness the existence of $0<\gamma$ so that
$0<\theta\leq\varepsilon^{\gamma}(\theta),\qquad\eta(\theta):=(\log|\log\varepsilon(\theta)|)^{-1}.$ (3)
We verify that $\theta^{\alpha}e^{\frac{C}{\eta^{2}}}$ is then bounded for any $\alpha>0$.
We now consider the standard paraxial envelope ansatz $u^{\theta}(z,x)=p\big{(}\frac{\eta z}{\varepsilon\theta},x\big{)}e^{-\frac{i\eta k_{0}z}{\varepsilon\theta^{2}}}$. After formally ignoring the backscattering term $(\frac{\varepsilon\theta}{\eta})^{2}\partial^{2}_{z}u^{\theta}$ in (1), we arrive at the paraxial approximation
$2ik_{0}\partial_{z}u^{\theta}+\frac{\eta}{\varepsilon}\Delta_{x}u^{\theta}+\frac{k_{0}^{2}}{(\eta\varepsilon\theta)^{\frac{1}{2}}}\nu\Big{(}\frac{\eta z}{\varepsilon\theta},x\Big{)}u^{\theta}=0,\quad z>0,x\in\mathbb{R}^{d}\,,$ (4)
with incident beam profile $u^{\theta}(0,x)=u^{\theta}_{0}(x)$. We will see below that the incident beams of our theory are broad, and more precisely functions of $\varepsilon^{\beta}x$ for $\beta\geq 1$. The justification of the paraxial regime from the full Helmholtz equation is difficult, with theoretical results obtained in [2, 15]. The envelope equation (4) is our starting point.
The objective of this paper is to show that in the limit $\theta\to 0$, $u^{\theta}$ satisfies in law a limiting model consistent with complex Gaussianity and in the diffusive regime with a scintillation index equal to $1$; see Corollary 3.5 below for notation. In this paper, we restrict ourselves to a setting of Gaussian fluctuations $\nu(z,x)$ to simplify the derivation.
The main step of our derivation is to show that the statistical moments of the paraxial solution $u^{\theta}$ in (4) are well approximated by the corresponding moments of the solution to the following Itô-Schrödinger equation:
$\mathrm{d}u^{\varepsilon}=\frac{i\eta}{2k_{0}\varepsilon}\Delta_{x}u^{\varepsilon}\mathrm{d}z-\frac{k_{0}^{2}R(0)}{8\eta^{2}}u^{\varepsilon}\mathrm{d}z+\frac{ik_{0}}{2\eta}u^{\varepsilon}\mathrm{d}B,\qquad u^{\varepsilon}(0,x)=u_{0}^{\varepsilon}(x).$ (5)
The above equation is deduced from (4) by formally replacing $\theta^{-\frac{1}{2}}\nu(\frac{z}{\theta},x)\mathrm{d}z$ by its central limit approximation $\mathrm{d}B(z,x)$, where $B(z,x)$ is a mean-zero Gaussian process satisfying $\mathbb{E}[B(z,x)B(z^{\prime},y)]=\min(z,z^{\prime})R(x-y)$ with spatial covariance $R(x)$ defined in (10) below. The product $u^{\varepsilon}\mathrm{d}B$ should be interpreted in a standard Itô form [10, 11].
That the complex circular Gaussian conjecture holds for $u^{\varepsilon}$ in the limit $\varepsilon\to 0$ was shown recently in [5]. Careful moment estimates of $u^{\theta}$ allow us to extend such a result to the paraxial model of wave propagation.
The rest of the paper is structured as follows. The next Section 2 defines the statistical moments of interest in the derivation and describes our hypotheses on the (Gaussian) random medium. Our main results are collected in Section 3. The proofs of the results are postponed to Section 4.
We restrict ourselves to the setting of deterministic incident beams (fully coherent sources). The generalization in [6] of the results in the Itô-Schrödinger regime of [5] to the case of partially coherence sources can also presumably be extended to the paraxial regime.
2 Description of the wavefield, random medium, and moments
For simplicity, we set $12k_{0}=$ 1 since this (rescaled) parameter does not have any qualitative influence, and assume that the amplitude $\nu$ is appropriately rescaled so that the paraxial equation becomes
$\partial_{z}u^{\theta}=\frac{i\eta}{\varepsilon}\Delta_{x}u^{\theta}+\frac{i}{(\eta\varepsilon\theta)^{\frac{1}{2}}}\nu\Big{(}\frac{\eta z}{\varepsilon\theta},x\Big{)}u^{\theta}.$ (6)
Define the (partial) Fourier transform of $u^{\theta}$ as
$\hat{u}^{\theta}(z,\xi)=\int\limits_{\mathbb{R}^{d}}u^{\theta}(z,x)e^{-i\xi\cdot x}\mathrm{d}x\,.$
From (6), we deduce that
$\partial_{z}\hat{u}^{\theta}=-\frac{i\eta}{\varepsilon}|\xi|^{2}\hat{u}^{\theta}+\frac{i}{(\eta\varepsilon\theta)^{\frac{1}{2}}}\int\limits_{\mathbb{R}^{d}}\hat{\nu}\Big{(}\frac{\eta z}{\varepsilon\theta},k\Big{)}\hat{u}^{\theta}(z,\xi-k)\frac{\mathrm{d}k}{(2\pi)^{d}},\quad\hat{u}^{\theta}(0,\xi)=\hat{u}^{\theta}_{0}(\xi)\,.$
For a collection of points $(X,Y)=(x_{1},\cdots,x_{p},y_{1},\cdots,y_{q})\in\mathbb{R}^{(p+q)d}$, define the $p+q$th moment of the wavefield $u^{\theta}$ at fixed $z$ by
$\mu_{p,q}^{\theta}(z,X,Y)=\mathbb{E}[\prod\limits_{j=1}^{p}u^{\theta}(z,x_{j})\prod\limits_{l=1}^{q}u^{\theta*}(z,y_{l})]\,.$ (7)
Similarly, define the corresponding (complex symmetrized) Fourier transform by
$\hat{\mu}_{p,q}^{\theta}(z,v)=\mathbb{E}[\prod\limits_{j=1}^{p}\hat{u}^{\theta}(z,\xi_{j})\prod\limits_{l=1}^{q}\hat{u}^{\theta*}(z,\zeta_{l})]\,,$
where $v=(\xi_{1},\cdots,\xi_{p},\zeta_{1},\cdots,\zeta_{q})$ denotes the complex symmetrized (i.e., with inner product $\sum x_{i}\cdot\xi_{i}-\sum y_{j}\cdot\zeta_{j}$) vector of variables dual to $(X,Y)$.
Gaussian random potential.
We assume that $\nu$ is a real valued Gaussian process stationary in $x$ and $z$ with covariance function
$\mathbb{E}[\nu(z,x)\nu(z^{\prime},x^{\prime})]=\mathfrak{C}(z-z^{\prime},x-x^{\prime})\,.$ (8)
In particular, this implies that the Fourier transform of the covariance function with respect to $x$ and $x^{\prime}$ is given by
$(2\pi)^{d}\hat{\mathfrak{C}}(z-z^{\prime},k)\delta(k+k^{\prime})\,.$
We assume that ^$\hat{\mathfrak{C}}(z,\cdot)$ is integrable in $z$, smooth in $k$ and rapidly decaying for large values of $z$. For concreteness and to simplify some derivations, we assume that for some $\tau_{R}>0$ and $\langle k\rangle:=\sqrt{1+|k|^{2}}$, we have
$\int_{\mathbb{R}^{d}}\langle{k}\rangle^{{d_{0}+3}}\hat{\mathfrak{C}}(z,k)\mathrm{d}k\leq\lambda_{1}(d_{0})e^{-\tau_{R}|z|},$ (9)
where we assume that $d_{0}\geq d$. Except for improved regularity results obtained when $d_{0}\geq d$ in Lemma 4.7 (and Lemma 4.2) and for better stochastic continuity properties in Theorem 3.3, there is no loss in generality to assume that $d_{0}=d$, the lateral spatial dimension. We assume the standard symmetries for correlation functions: $\mathfrak{C}(z,x)=\mathfrak{C}(-z,x)=\mathfrak{C}(z,-x)$ and denote by $R(x)$ the lateral spatial covariance
$R(x)=\int\limits_{-\infty}^{\infty}\mathfrak{C}(z,x)\mathrm{d}z\,.$ (10)
We next assume that $R\in{\mathsf{L}}^{1}(\mathbb{R}^{d})\cap{\mathsf{L}}^{\infty}(\mathbb{R}^{d})$ with the Fourier transform $\hat{R}(k)=\hat{R}(-k)\leq\hat{{\mathsf{R}}}(k)=\hat{{\mathsf{R}}}(|k|)$ for some radially symmetric function $\hat{{\mathsf{R}}}$.
We deduce from (9) that $\langle k\rangle^{2}\hat{R}(k)\in{\mathsf{L}}^{1}(\mathbb{R}^{d})$, and when $d\geq 3$ further assume that $\langle k\rangle^{d-2}\hat{R}(k)\in{\mathsf{L}}^{\infty}(\mathbb{R}^{d})$. We finally assume that $R(x)$ is maximal at $x=0$ and that $\Xi=\nabla^{2}R(0)$ exists and is (strictly) negative definite.
As in [11], we denote the potential in the rescaled coordinates by
$\nu^{\theta}_{z}(\cdot)=\nu\Big{(}\frac{\eta z}{\varepsilon\theta},\cdot\Big{)}\,,$ (11)
and the corresponding covariance as
${\mathfrak{C}}^{\theta}(s,x)={\mathfrak{C}}\Big{(}\frac{\eta s}{\varepsilon\theta},x\Big{)},$ (12)
with the Fourier transform ^$\hat{\mathfrak{C}}^{\theta}(s,k)$ defined in a similar manner.
Assumptions on the source.
We assume a deterministic incident beam profile of the form
$u_{0}^{\theta}(x)=u_{0}(\varepsilon^{\beta}x)\,,$ (13)
with $u_{0}\in\mathcal{S}(\mathbb{R}^{d})$ and $\beta\geq 1$. This profile corresponds to a wide beam when $\varepsilon\ll 1$. The scaling $\beta=1$ is arguably the more interesting one. The case $\beta>1$ is a good model for an incident plane wave. Note that in the Fourier domain,
$\hat{u}^{\theta}_{0}(k)=\varepsilon^{-\beta d}\hat{u}_{0}\Big{(}\frac{k}{\varepsilon^{\beta}}\Big{)}\,.$
Due to this scaling, we have that for all $n\geq 0$,
$\|\langle k\rangle^{n}\hat{u}^{\theta}_{0}(k)\|\leq\|\langle\varepsilon^{-\beta}k\rangle^{n}\hat{u}^{\theta}_{0}(k)\|=\|\langle k\rangle^{n}\hat{u}_{0}(k)\|\leq C\,$
uniformly in $\varepsilon\in(0,1]$.
[…]
Spatially rescaled random vector.
As in [5], we define the spatially rescaled random process $\phi^{\theta}$ as
$\phi^{\theta}(z,r,x)=u^{\theta}(z,\varepsilon^{-\beta}r+\eta x).$ (14)
It is for such a rescaled wavefield that we can expect a non-trivial limit as $\theta\to 0$ and $\varepsilon\to 0$.
An element of the convergence result is the identification of finite dimensional distributions of $\phi^{\theta}$. Thus for $X=(x_{1},\cdots,x_{N})\in\mathbb{R}^{Nd}$, we also define a vector of such processes as
$\Phi^{\theta}(z,r,X)=(\phi^{\theta}(z,r,x_{1}),\cdots,\phi^{\theta}(z,r,x_{N}))\,.$ (15)
The moments of this random vector will be shown to be well approximated by that of a circularly symmetric Gaussian random vector in the limit $(\theta,\varepsilon)\to 0$ (provided (3) holds).
3 Main results
The limiting approximations for $\psi^{\theta}$ as $\theta\to 0$ in the kinetic and diffusive regimes are the same as for the Itô-Schrödinger model (as $\varepsilon\to 0$) treated in [5]. The limits are described by second-order moments, which we now recall.
In the kinetic regime where $\eta=1$, we define the second-order moment $M_{1,1}(z,r,x,y)$ as
$M_{1,1}(z,r,x,y)=\begin{cases}|u_{0}(r)|^{2}\mathcal{Q}(y-x,0),\quad&\beta>1\\ \displaystyle\int\limits_{\mathbb{R}^{d}}e^{i\xi\cdot(r-r^{\prime})}|u_{0}(r^{\prime})|^{2}\mathcal{Q}(y-x,\xi)\frac{\mathrm{d}\xi\mathrm{d}r^{\prime}}{(2\pi)^{d}},\quad&\beta=1\,,\end{cases}$ (16)
where
$\mathcal{Q}(\tau,\zeta)=\exp\Big{(}\frac{z}{\eta^{2}}\int\limits_{0}^{1}[R\big{(}\tau+2sz\zeta\big{)}-R(0)]\mathrm{d}s\Big{)}\,.$ (17)
In the diffusive regime, characterized by $\eta=(\log|\log\varepsilon|)^{-1}$, we define $M_{1,1}(z,r,x,y)$ as
$M_{1,1}(z,r,x,y)=\begin{cases}|u_{0}(r)|^{2}e^{\frac{z}{2}(y-x)^{\mathrm{t}}\Xi(y-x)},\quad&\beta>1\\[5.69054pt] e^{\frac{1}{8}z(y-x)^{\mathrm{t}}\Xi(y-x)}e^{-\frac{3i}{4z}(y-x)\cdot r}G(z^{3},r)\ast[e^{\frac{3i}{4z}(y-x)\cdot r}|u_{0}(r)|^{2}],\quad&\beta=1\,,\end{cases}$ (18)
where $\Xi=\nabla^{2}R(0)$ (is negative definite) and $G(t,r)$ is the Green’s function to the evolution equation
$\partial_{t}G+\frac{2}{3}\nabla_{r}\cdot(\Xi\nabla_{r})G(t,r)=0,\quad G(0,r)=\delta_{0}(r)\,.$ (19)
We are now ready to state our main theorems. We assume (3) throughout and recall that $\Phi^{\theta}$ is defined in (15).
Theorem 3.1.
In the kinetic regime $\eta=1$, $\Phi^{\theta}(z,r,X)\Rightarrow\Phi(z,r,X)$ in distribution as $\theta\to 0$ where $\Phi(z,r,X)$ is a random vector with elements $\{\phi_{j}\}_{j=1}^{N}$ satisfying $\mathbb{E}\phi_{j}=M_{1,0}(z,r,x_{j})$ and $Z=\Phi-\mathbb{E}\Phi$ a mean zero circular complex Gaussian random vector with elements $\{Z_{j}\}_{j=1}^{N}$ satisfying
$\mathbb{E}[Z_{j}Z^{\ast}_{l}]=M_{1,1}(z,r,x_{j},x_{l})-M_{1,0}(z,r,x_{j})M_{0,1}(z,r,x_{l}),\quad\mathbb{E}[Z_{j}Z_{l}]=0\,.$ (20)
Here, $M_{0,1}(z,r,x)=M_{1,0}^{\ast}(z,r,x)=e^{-\frac{R(0)z}{2}}u_{0}(r)$ and $M_{1,1}(z,r,x,y)$ is given by (16).
Theorem 3.2.
In the diffusive regime $\eta=(\log|\log\varepsilon|)^{-1}$, $\Phi^{\theta}(z,r,X)\Rightarrow\Phi(z,r,X)$ in distribution as $\theta\to 0$ where $\Phi(z,r,X)$ is a mean zero circular complex Gaussian random vector with elements $\{\phi_{j}\}_{j=1}^{N}$ satisfying
$\mathbb{E}[\phi_{j}\phi_{l}^{\ast}]=M_{1,1}(z,r,x_{j},x_{l}),\quad\mathbb{E}[\phi_{j}\phi_{l}]=0\,,$ (21)
with $M_{1,1}(z,r,x,y)$ given by (18).
As in [5], the results in Theorems 3.1 and 3.2 at $(z,r,X)$ fixed characterize finite dimensional distributions of the process $x\to\phi^{\theta}(z,r,x)$ defined in (14) at $(z,r)$ fixed. The proof of convergence of the whole process follows from a tightness and stochastic continuity result of independent interest, which we now state.
Theorem 3.3.
$\sup\limits_{0\leq s\leq z}\mathbb{E}|\phi^{\theta}(s,r,x+h)-\phi^{\theta}(s,r,x)|^{2n}\leq C(n,\alpha,z)|h|^{2n\alpha}$ (22)
for $\|h\|_{2}<1$ uniformly in $\theta\in(0,\frac{1}{2}]$. Here, $\alpha\in(0,1]$ in the kinetic regime and $\alpha\in(0,1)$ in the diffusive regime. Let $n$ be such that $2n\alpha\geq d+2n\alpha_{-}$, $\alpha_{-}\in(0,\alpha)$. Then the process $\phi^{\theta}(z,r)$ is tight on $C^{0,\alpha_{-}}(\mathbb{R}^{d})$ and there exists a Hölder continuous version of $\phi^{\theta}(z,r)$ on $C^{0,\alpha_{-}}(\mathbb{R}^{d})$.
All results are proved in Section 4.
As in [5, Corollary 2.8], we deduce the following result of convergence in distribution:
Corollary 3.4 (Convergence of processes).
For $z$ and $r$ fixed and under the hypotheses of Theorem 3.3, the process $\phi^{\theta}(z,r,\cdot)$ converges in distribution as probability measures on $C^{0,\alpha_{-}}(\mathbb{R}^{d})$ to its unique limit, which is characterized by limits of finite dimensional distributions shown to converge in Theorems 3.1 and 3.2 in the kinetic and diffusive regimes.
For completeness, we recall the following corollary on the scintillation index. This is [5, Corollary 2.5], written with $\beta=1$ to simplify the presentation, and to which we refer for derivation and details.
Corollary 3.5 (Scintillation in the diffusive regime).
Assume $\beta=1$ and define the intensity
$I^{\theta}(z,r,x)=|u^{\theta}(z,\varepsilon^{-1}r+\eta x)|^{2}\,.$ (23)
Then (at fixed $(z,r)$) $I^{\theta}\boldsymbol{\Rightarrow}I$ in distribution as $\theta\to 0$ where $I(z,r)$ is distributed according to an exponential law with $\mathbb{E}[I](z,r)=(G(z^{3},\cdot)*|u_{0}|^{2})(z,r)$, where $G$ is the fundamental solution to the diffusion equation (19) and * denotes convolution (in the $r-$variable). In particular, the scintillation index
${\mathsf{S}}(z,r):=\frac{\mathbb{E}[I^{2}]-\mathbb{E}[I]^{2}}{\mathbb{E}[I]^{2}}=1,\quad\forall z>0,\ r\in\mathbb{R}^{d}\,.$ (24)
The above corollary fully characterizes the scintillation regime with a scintillation index uniformly equal to $1$ corresponding to a fully developed speckle pattern. The above diffusion with kernel $G(z^{3},\cdot)$ is naturally written in the ‘time’ variable $z^{3}$ and lateral ‘spatial’ variables $r$, resulting in a very anomalous diffusion regime with $r\sim z^{\frac{3}{2}}$ for large $z$ reflecting similar beam spreading in the paraxial and Itô-Schrödinger regimes.
Scaling in applications.
We briefly revisit the scaling (2) and its practical relevance. There, $x$ was measured in terms of the smallest scale of turbulence, let us call it $l_{0}$, typically taken in practice to be in the $mm$ range. This means that medium fluctuations should be understood as $\nu=\nu\big{(}\frac{z}{l_{0}},\frac{x}{l_{0}}\big{)}$. With $Z$ a dimension-less parameter characterizing longitudinal distance of propagation, $w_{0}$ a parameter describing the width of the source, and $\sigma$ a covariance parameter for the random fluctuations $\nu=\sigma\tilde{\nu}$, we have the scaling and paraxial model after neglecting backscattering:
$\frac{z}{l_{0}Z}\to z,\quad\frac{x}{l_{0}}\to x,\qquad 2i\Big{(}\frac{k_{0}l_{0}}{Z}\Big{)}\partial_{z}u+\Delta_{x}u+k_{0}^{2}l_{0}^{2}\sigma\tilde{\nu}(zZ,x)u=0,\quad u(0,x)=u_{0}\Big{(}\frac{l_{0}x}{w_{0}}\Big{)}.$
Consider the scaling with $\beta=1$ for concreteness. The parameters $(\theta,\varepsilon,\eta)$ are then defined as
$\theta=\frac{1}{k_{0}l_{0}},\quad\varepsilon=\frac{l_{0}}{w_{0}},\quad\eta=\frac{\varepsilon Z}{k_{0}l_{0}}=\frac{Z}{k_{0}w_{0}}.$
We then verify that the above paraxial model gives (4) when $\sigma^{2}Z^{3}l_{0}^{2}=w_{0}^{2}$. This is equivalent to $(\sigma^{2}l_{0}^{2}k_{0}^{2})\eta^{2}Z=1$.
A typical wavelength used in practice is $\lambda\approx 10^{-6}m$ and corresponds to $k_{0}=10^{6}m^{-1}$. Under weak turbulence assumptions [9, 11], the variations in the refractive index are consistent with $\sigma=10^{-4}$ with a fluctuation lengthscale $l_{0}=10^{-3}m$. Thus, $\theta=10^{-3}$ and $\sigma^{2}l_{0}^{2}k_{0}^{2}=10^{-2}$.
We thus observe that we reach the kinetic regime $\eta=1$ for a propagation distance of order $Zm=10^{2}m$. We are in the scintillation regime when $\varepsilon\ll 1$ or $l_{0}\ll w_{0}$, which is the case for the beam widths of a (quite a) few millimeters to meters found in applications [25, 28, 27, 1].
The diffusive regime is reached for $\eta\ll 1$, corresponding to $Z\gg 10^{2}$, with values of $Zm$ in the range $10^{3}-10^{4}m$ typical in applications. For $Z=10^{4}$ so that $\eta=10^{-1}$ for concreteness, the proposed theory has a domain of validity for beam width $w_{0}=10^{-3}/\varepsilon\gg 10^{-3}/\eta=10^{-2}m$, i.e., for beams with a diameter at least in the tens of centimeters to meter scale.
4 Proofs of the main results
The main strategy of the proof of the above results is to show that the exact moment equations available in the Itô-Schrödinger regime are approximately valid in the paraxial regime. Once such a result is established, the characterization of the limits of finite dimensional distributions of $\phi^{\theta}$ closely follows the derivation in [5].
4.1 Phase compensated wave field and preparatory calculations
As in [4, 22] and [5], approximations to the wavefield are best performed in the Fourier domain and after ‘removing’ or ‘compensating for’ the effect of free propagation in a homogeneous medium. The phase compensated wave field $\tilde{u}^{\theta}(z,\xi)=\hat{u}^{\theta}(z,\xi)e^{\frac{i\eta z}{\varepsilon}|\xi|^{2}}$ solves the evolution equation
$\partial_{z}\tilde{u}^{\theta}=\frac{i}{(\eta\varepsilon\theta)^{\frac{1}{2}}}\int\limits_{\mathbb{R}^{d}}\hat{\nu}\Big{(}\frac{\eta z}{\varepsilon\theta},k\Big{)}\tilde{u}^{\theta}(z,\xi-k)e^{\frac{i\eta z}{\varepsilon}g(\xi,k)}\frac{\mathrm{d}k}{(2\pi)^{d}},\quad\tilde{u}^{\theta}(0,\xi)=\hat{u}^{\theta}_{0}(\xi)\,,$ (25)
where $g(\xi,k)=|\xi|^{2}-|\xi-k|^{2}=-|k|^{2}+2k\cdot\xi$. In terms of the phase compensated wave field, the $p+q$th moment $\hat{\mu}^{\theta}_{p,q}$ is then given by
$\hat{\mu}_{p,q}^{\theta}(z,v)=\Pi_{p,q}^{\theta}(z,v)\mathbb{E}[\psi^{\theta}_{p,q}(z,v)],$
where
$\Pi_{p,q}^{\theta}(z,v)=e^{-\frac{i\eta z}{\varepsilon}(\sum_{j=1}^{p}|\xi_{j}|^{2}-\sum_{l=1}^{q}|\zeta_{l}|^{2})},$
and $\psi^{\theta}_{p,q}$ is a product of $p+q$ copies of the phase compensated wave field given by
$\psi^{\theta}_{p,q}(z,v)=\prod\limits_{j=1}^{p}\tilde{u}^{\theta}(z,\xi_{j})\prod\limits_{l=1}^{q}\tilde{u}^{\theta*}(z,\zeta_{l}).$
Denote by $\Psi^{\theta}_{p,q}$ the $p+q$th moment of the phase compensated field:
$\Psi_{p,q}^{\theta}(z,v)=\mathbb{E}\psi_{p,q}^{\theta}(z,v).$ (26)
We drop the $p,q$ dependence on $\psi$ for notational convenience. From (25), we have
$\partial_{z}\psi^{\theta}=\frac{i}{(\eta\varepsilon\theta)^{\frac{1}{2}}}L^{\theta}(z,\nu^{\theta}_{z})\psi^{\theta},\quad\psi^{\theta}(0,v)=\prod\limits_{j=1}^{p}\hat{u}^{\theta}_{0}(\xi_{j})\prod\limits_{l=1}^{q}\hat{u}^{\theta*}_{0}(\zeta_{l})\,,$ (27)
where we recall that $\nu^{\theta}_{z}(\cdot)=\nu\big{(}\frac{\eta z}{\varepsilon\theta},\cdot\big{)}$ and the operator $L^{\theta}$ is given by
$L^{\theta}(s,\nu_{z})\psi=\int\limits_{\mathbb{R}^{d}}\hat{\nu}(z,k)\Big{[}\sum\limits_{j=1}^{p}e^{\frac{i\eta s}{\varepsilon}g(\xi_{j},k)}\psi(\xi_{j}-k)-\sum\limits_{l=1}^{q}e^{-\frac{i\eta s}{\varepsilon}g(\zeta_{l},-k)}\psi(\zeta_{l}+k)\Big{]}\frac{\mathrm{d}k}{(2\pi)^{d}}\,,$ (28)
where we explicitly indicate only the variables in $\psi$ that are shifted by ±$\pm k$.
Lemma 4.1.
$\psi^{\theta}$ can be expanded as
$\psi^{\theta}(z,v)=\sum\limits_{n\geq 0}\Big{(}\frac{i}{(\eta\varepsilon\theta)^{\frac{1}{2}}}\Big{)}^{n}\psi^{\theta}_{n}(z,v)\,,$ (29)
where $\psi^{\theta}_{0}(z,v)=\psi^{\theta}(0,v)$ and for $n\geq 1$,
$\psi^{\theta}_{n}(z,v)=\int\limits_{[0,z]^{n}_{<}}\Big{(}\prod\limits_{j=1}^{n}L^{\theta}(s_{j},\nu^{\theta}_{s_{j}})\Big{)}\psi^{\theta}(0,v)\prod\limits_{j=1}^{n}\mathrm{d}s_{j},$
where $[0,z]^{n}_{<}$ denotes the simplex $0\leq s_{1}\leq\cdots\leq s_{n}\leq z$. Moreover,
$\mathbb{E}\psi^{\theta}_{2N}(z,v)=\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{Nd}}\sum\limits_{P^{2N}}\sum\limits_{j=1}^{(p+q)^{2N}}e^{\frac{i\eta}{\varepsilon}G^{\prime}_{j}(\vec{s},\vec{k},v)}\psi^{\theta}(0,v-A^{\prime}_{j}\vec{k})\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$ (30)
Here, $A_{j}^{\prime}$ is a $(p+q)d\times Nd$ block diagonal matrix with uniformly bounded entries and $G_{j}^{\prime}$ is real valued $\forall j$.
Proof.
Using (27) and a standard Duhamel expansion, we arrive at (29). Defining these iterated integrals and the convergence in (29) in a square integrable sense, can be justified as in, e.g., the proof of [4, Proposition 2.1]. Using the definition of $L^{\theta}$, this can be further expanded as
$\psi^{\theta}_{n}(z,v)=\int\limits_{[0,z]^{n}_{<}}\int\limits_{\mathbb{R}^{nd}}\Big{(}\prod\limits_{j=1}^{n}\hat{\nu}^{\theta}_{s_{j}}(k_{j})\Big{)}\sum\limits_{j=1}^{(p+q)^{n}}e^{\frac{i\eta}{\varepsilon}G_{j}(\vec{s},\vec{k},v)}\psi^{\theta}(0,v-A_{j}\vec{k})\prod\limits_{j=1}^{n}\frac{\mathrm{d}k_{j}\mathrm{d}s_{j}}{(2\pi)^{d}}\,,$ (31)
where $\vec{s}=(s_{1},\cdots,s_{n})$ and $\vec{k}=(k_{1},\cdots,k_{n})$. Here, $A_{j}$ are matrices and $G_{j}$ are real valued functions which we now define.
$A_{j}$ are $(p+q)d\times nd$ block matrices with $d\times d$ blocks such that $A_{j}$ has exactly one non zero entry per column block. This non zero entry equals 1 when it appears in the first $pd$ rows and equals $-1$ otherwise. $G_{j}$ can be defined recursively as follows:
Let $\epsilon_{m}=1$ if $m=1,\cdots,p$ and equal to $-1$ if $m=p+1,\cdots,p+q$. Also, let
$G_{1,m}(\vec{s},\vec{k},v)=s_{1}\epsilon_{m}g(v_{m},\epsilon_{m}k_{1})\,.$
Now, for $r\geq 1$, let $G_{r+1,j}(\vec{s},\vec{k},v)$ be given by
$G_{r+1,(p+q)(m^{\prime}-1)+m}(\vec{s},\vec{k},v)$ $=s_{r+1}\epsilon_{m}g(v_{m},\epsilon_{m}k_{r+1})$
$+G_{r,m^{\prime}}(s_{1},\cdots,s_{r},k_{1},\cdots,k_{r},v_{1},\cdots,v_{m}-\epsilon_{m}k_{r+1},v_{m+1},\cdots,v_{p+q})\,,$
where $m^{\prime}=1,\cdots,(p+q)^{r}$ and $m=1,\cdots,p+q$. Here, for $r$ fixed, $j$ varies from $1$ to $(p+q)^{r}$. Finally set $G_{j}$ to be $G_{n,j}$.
From the Gaussian assumption on $\nu$, chaos with odd $n$ vanish after taking expectation. Applying Wick’s theorem [23], we have
$\mathbb{E}\psi^{\theta}_{2N}(z,v)=\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{2Nd}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}(2\pi)^{d}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\delta(k_{j^{\prime}}+k_{l^{\prime}})$
×$\times\sum\limits_{j=1}^{(p+q)^{2N}}e^{\frac{i\eta}{\varepsilon}G_{j}(\vec{s},\vec{k},v)}\psi^{\theta}(0,v-A_{j}\vec{k})\prod\limits_{j=1}^{2N}\frac{\mathrm{d}k_{j}\mathrm{d}s_{j}}{(2\pi)^{d}}\,.$
Here, $P^{2N}$ denotes a permutation of pairings of $2N$ indices. The Dirac measure reduces integrals over $\vec{k}\in\mathbb{R}^{2Nd}$ to an integral over $\mathbb{R}^{Nd}$. This gives (30).
In (30), $A_{j}^{\prime}$ are $(p+q)d\times Nd$ block diagonal matrices of $d\times d$ blocks with either no non zero entries or exactly two non zero entries (either 1 or -1) per column block. Also, for a fixed arrangement $P^{2N}$ with $(l,m)\in P^{2N}$, $G_{j}^{\prime}(\vec{s},\vec{k},v)=G_{j}(\vec{s},B_{P^{2N}}(\vec{k}),v)$ where $B_{P^{2N}}(\vec{k})$ is a block matrix of size $2Nd\times 1$ with $k_{1},\cdots,k_{N}$ appearing as $l$th entries and $-k_{1},-k_{2},\cdots,-k_{N}$ appearing as corresponding $m$th entries. For example, when $N=2$, there are three possibilities for $P^{4}$: if $P^{4}=\{(1,2),(3,4)\}$, $B_{P^{4}}(\vec{k})=(k_{1},-k_{1},k_{2},-k_{2})^{\top}$ and if $P^{4}=\{(1,3),(2,4)\}$, $B_{P^{4}}(\vec{k})=(k_{1},k_{2},-k_{1},-k_{2})^{\top}$ and finally if $P^{4}=\{(1,4),(2,3)\}$, $B_{P^{4}}(\vec{k})=(k_{1},k_{2},-k_{2},-k_{1})^{\top}$. ∎
The analysis of $\Psi^{\theta}_{p,q}$ in (26) could also be carried out by computing products of Duhamel expansions of $\psi^{\theta}$ (with $p=1$ and $q=0$). It proved more convenient to write a Duhamel expansion directly for $\psi^{\theta}_{p,q}$ in (29) as it is more closely related to the moment equations available for the Itô-Schrödinger solution and used in the analysis in [5]. For the rest of the paper, all we need of the above computations is the form of the moments in (30) and the fact that the entries of all matrices $A^{\prime}_{j}$ are bounded and that the phase terms $G_{j}^{\prime}$ are real-valued. This allows us to obtain the following moment estimates.
Lemma 4.2.
We have the following estimates:
$\|\mathbb{E}\psi^{\theta}(z)\|\leq\|\psi^{\theta}(0)\|e^{\frac{(p+q)^{2}R(0)z}{2\eta^{2}}}\,,$ (32)
and
$\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\langle v_{n}\rangle^{\alpha^{\prime}}\langle v_{n^{\prime}}\rangle^{\alpha^{\prime}}\Big{(}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\Big{)}\mathbb{E}\psi^{\theta}(z)\|\leq\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha+\alpha^{\prime}}\psi^{\theta}(0)\|e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\,,$ (33)
with $\lambda$ defined in (36) below for $\alpha^{\prime}\in[0,1]$ and $0\leq\alpha\leq\frac{d_{0}+1}{p+q}$.
Proof.
From the chaos expansion (29), (30), we have
$\|\mathbb{E}\psi^{\theta}(z)\|\leq\sum\limits_{N\geq 0}\frac{1}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{Nd}}\sum\limits_{P^{2N}}\sum\limits_{j=1}^{(p+q)^{2N}}\|\psi^{\theta}(0,v-A_{j}^{\prime}\vec{k})\|\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
Since a shift in $v$ does not affect the total variation norm, we have that
$\|\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{Nd}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}$
$=\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}_{<}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},0)\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
Now we can symmetrize the simplex:
$\|\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{1}{(2N)!}\frac{(p+q)^{2N}}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},0)\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}$
$=\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{2^{N}N!}\Big{(}\frac{1}{\eta\varepsilon\theta}\int\limits_{[0,z]^{2}}\mathfrak{C}^{\theta}(s_{1}-s_{2},0)\mathrm{d}s_{1}\mathrm{d}s_{2}\Big{)}^{N}\,.$
Here we have used the fact that $|P^{2N}|=\frac{(2N)!}{2^{N}N!}$. From a change of variables $\frac{\eta}{\varepsilon\theta}(s_{1}-s_{2})\to s_{2}$, we have
$\|\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{(p+q)^{2N}}{2^{N}N!}\Big{(}\frac{z}{\eta^{2}}R(0)\Big{)}^{N}=\|\psi^{\theta}(0)\|e^{\frac{(p+q)^{2}R(0)z}{2\eta^{2}}}\,.$
For the second part, we note that
$\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha_{m}}\mathbb{E}\psi^{\theta}(z)\|$ $\leq\sum\limits_{N\geq 0}\frac{1}{(\eta\varepsilon\theta)^{N}}\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{Nd}}\sum\limits_{P^{2N}}\sum\limits_{j=1}^{(p+q)^{2N}}\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha_{m}}\psi^{\theta}(0,v-A_{j}^{\prime}\vec{k})\|$
×$\times\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,,$
where $\alpha_{m}=\alpha$ if $m\neq n,n^{\prime}$ and $\alpha_{m}=\alpha^{\prime}+\alpha$ if $m=n,n^{\prime}$.
For each arrangement $A_{j}^{\prime}$, we make a change of variables $v_{m}-[A^{\prime}_{j}\vec{k}]_{m}$ so that
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha_{m}}\psi^{\theta}(0,v-A_{j}^{\prime}\vec{k})\|$ $=\|\prod\limits_{m=1}^{p+q}\langle v_{m}+[A^{\prime}_{j}\vec{k}]_{m}\rangle^{\alpha_{m}}\psi^{\theta}(0,v)\|\,.$ (34)
We note that
$1\leq\langle\xi+k\rangle^{2}=1+|\xi+k|^{2}\leq 1+2(|\xi|^{2}+|k|^{2})\leq 2\langle\xi\rangle^{2}\langle k\rangle^{2}\,.$
Using Lemma 4.1, this can be generalized to give
$1\leq\langle v_{m}+[A^{\prime}_{j}\vec{k}]_{m}\rangle^{\alpha_{m}}\leq C^{N}\langle v_{m}\rangle^{\alpha_{m}}\prod\limits_{j=1}^{N}\langle[\vec{k}]_{j}\rangle^{\alpha_{m}}\,.$ (35)
This gives
$\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha_{m}}\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha^{\prime}+\alpha}\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{C^{(p+q)N}(p+q)^{2N}}{(\eta\varepsilon\theta)^{N}}$
×$\times\int\limits_{[0,z]^{2N}_{<}}\int\limits_{\mathbb{R}^{Nd}}\sum\limits_{P^{2N}}\prod\limits_{(j^{\prime},l^{\prime})\in P^{2N}}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\langle k_{j^{\prime}}\rangle^{2+(p+q)\alpha}\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
Note that
$\int\limits_{\mathbb{R}}\int\limits_{\mathbb{R}^{d}}\langle k\rangle^{n}\hat{\mathfrak{C}}(s,k)\frac{\mathrm{d}k\mathrm{d}s}{(2\pi)^{d}}=\|\langle k\rangle^{n}\hat{R}(k)\|\,.$
Then as before, symmetrizing the simplex and making an appropriate change of variables gives
$\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha_{m}}\mathbb{E}\psi^{\theta}(z)\|$ $\leq\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha^{\prime}+\alpha}\psi^{\theta}(0)\|\sum\limits_{N\geq 0}\frac{C^{(p+q)N}(p+q)^{2N}}{2^{N}N!\eta^{2N}}z^{N}\|\langle k\rangle^{2+(p+q)\alpha}\hat{R}(k)\|^{N}$
$\leq\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha^{\prime}+\alpha}\psi^{\theta}(0)\|e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}},$
with
$\lambda[R,p,q,z,\alpha]=\frac{1}{2}C^{p+q}(p+q)^{2}z\|\langle k\rangle^{2+(p+q)\alpha}\hat{R}(k)\|\,.$ (36)
4.2 Convergence of moments: from paraxial to Itô-Schrödinger
The main results in Theorems 3.1 and 3.2 rely on the following decomposition, which is the main result of this paper:
Theorem 4.3.
The moments of the phase compensated wave field in (26) may be decomposed as
$\Psi_{p,q}^{\theta}(z,v)=\mathcal{B}^{\theta}_{p,q}(z,v)+\mathscr{R}^{\theta}_{p,q}(z,v)\,,$
where $\mathcal{B}^{\theta}_{p,q}(z,v)$ is the solution to
$\mathcal{B}^{\theta}_{p,q}(z,v)=\psi^{\theta}(0,v)+\int\limits_{0}^{z}\mathcal{L}^{\theta}(s)\mathcal{B}^{\theta}_{p,q}(s)\mathrm{d}s\,,$ (37)
with $\mathcal{L}^{\theta}$ given by
$\mathcal{L}^{\theta}(z)\psi$ $=\frac{1}{\eta^{2}}\int\limits_{\mathbb{R}^{d}}\hat{R}(k)\Big{[}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}e^{\frac{i\eta z}{\varepsilon}[g(\xi_{j},k)-g(\zeta_{l},k)]}\psi(\xi_{j}-k,\zeta_{l}-k)$ (38)
$-\sum\limits_{1\leq j<j^{\prime}\leq p}e^{\frac{i\eta z}{\varepsilon}[g(\xi_{j},k)+g(\xi_{j^{\prime}},-k)]}\psi(\xi_{j}-k,\xi_{j^{\prime}}+k)$
$-\sum\limits_{1\leq l<l^{\prime}\leq q}e^{-\frac{i\eta z}{\varepsilon}[g(\zeta_{l},k)+g(\zeta_{l^{\prime}},-k)]}\psi(\zeta_{l}-k,\zeta_{l^{\prime}}+k)-\frac{p+q}{2}\psi\Big{]}\frac{\mathrm{d}k}{(2\pi)^{d}}\,,$
and $\mathscr{R}^{\theta}_{p,q}(z,v)$ satisfying
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{p,q}(z,v)\|\leq c(p,q,z)\frac{\theta}{\eta^{2}}e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{2}\psi^{\theta}(0,v)\|\,,$ (39)
where $\lambda[R,p,q,z,\alpha]=\frac{1}{2}C^{p+q}(p+q)^{2}z\|\langle k\rangle^{2+(p+q)\alpha}\hat{R}(k)\|$ for a fixed constant $C$ and $0\leq\alpha\leq\frac{d_{0}+1}{p+q}$.
Proof.
Iterating (27) twice in integral form gives
$\mathbb{E}\psi^{\theta}(z,v)=\mathbb{E}\psi^{\theta}(0,v)-\frac{1}{\eta\varepsilon\theta}\int\limits_{0}^{z}\int\limits_{0}^{s}\mathbb{E}[L^{\theta}(s,\nu^{\theta}_{s})L^{\theta}(s^{\prime},\nu^{\theta}_{s^{\prime}})\psi^{\theta}(s^{\prime})]\mathrm{d}s^{\prime}\mathrm{d}s\,.$ (40)
We use the following relations in turn
$\psi^{\theta}$ $=\mathbb{E}\psi^{\theta}+(\psi^{\theta}-\mathbb{E}\psi^{\theta})$
$L^{\theta}(s^{\prime},\nu^{\theta}_{s^{\prime}})$ $=L^{\theta}(s,\nu^{\theta}_{s^{\prime}})+(L^{\theta}(s^{\prime},\nu^{\theta}_{s^{\prime}})-L^{\theta}(s,\nu^{\theta}_{s^{\prime}}))$
[…]
to obtain that
$\mathbb{E}\psi^{\theta}(z,v)=\mathbb{E}\psi^{\theta}(0,v)+\int\limits_{0}^{z}\mathcal{L}^{\theta}(s)\mathbb{E}\psi^{\theta}(s)\mathrm{d}s+\mathscr{R}^{\theta}_{p,q}(z,v)\,,$
where the leading operator $\mathcal{L}^{\theta}(s)$ and the residual $\mathscr{R}^{\theta}_{p,q}(z,v)$ are given by
$\mathcal{L}^{\theta}(s)$ $=$ $-\dfrac{1}{\eta\varepsilon\theta}\displaystyle\int\limits_{s}^{\infty}\mathbb{E}[L^{\theta}(s,\nu^{\theta}_{s})L^{\theta}(s,\nu^{\theta}_{s^{\prime}})]\mathrm{d}s^{\prime}$ (41)
$\mathscr{R}^{\theta}_{p,q}(z,v)$ $=$ $\mathscr{R}^{\theta}_{1}(z,v)+\mathscr{R}^{\theta}_{2}(z,v)+\mathscr{R}^{\theta}_{3}(z,v),$ (42)
where
$\mathscr{R}^{\theta}_{1}(z,v)$ $=-\frac{1}{\eta\varepsilon\theta}\int\limits_{0}^{z}\int\limits_{0}^{s}\Big{(}\mathbb{E}[L^{\theta}(s,\nu^{\theta}_{s})L^{\theta}(s^{\prime},\nu^{\theta}_{s^{\prime}})\psi^{\theta}(s^{\prime})]-\mathbb{E}[L^{\theta}(s,\nu^{\theta}_{s})L^{\theta}(s^{\prime},\nu^{\theta}_{s^{\prime}})]\mathbb{E}\psi^{\theta}(s^{\prime})\Big{)}\mathrm{d}s^{\prime}\mathrm{d}s$ (43)
$\mathscr{R}^{\theta}_{2}(z,v)$ $=\frac{1}{\eta\varepsilon\theta}\int\limits_{0}^{z}\int\limits_{0}^{s}\big{(}\mathbb{E}[L^{\theta}(s^{\prime},\nu^{\theta}_{s^{\prime}})L^{\theta}(s^{\prime},\nu^{\theta}_{s})]-\mathbb{E}[L^{\theta}(s,\nu^{\theta}_{s})L^{\theta}(s^{\prime},\nu^{\theta}_{s^{\prime}})]\big{)}\mathbb{E}\psi^{\theta}(s^{\prime})\mathrm{d}s^{\prime}\mathrm{d}s$
$\mathscr{R}^{\theta}_{3}(z,v)$ $=\frac{1}{\eta\varepsilon\theta}\int\limits_{z}^{\infty}\int\limits_{0}^{z}\mathbb{E}[L^{\theta}(s^{\prime},\nu^{\theta}_{s^{\prime}})L^{\theta}(s^{\prime},\nu^{\theta}_{s})]\mathbb{E}\psi^{\theta}(s^{\prime})\mathrm{d}s^{\prime}\mathrm{d}s\,.$
From (28) we observe that
$\mathbb{E}[L^{\theta}(z_{1},\nu^{\theta}_{s_{1}})L^{\theta}(z_{2},\nu^{\theta}_{s_{2}})]\psi$
$=-\int\limits_{\mathbb{R}^{2d}}\mathbb{E}[\hat{\nu}^{\theta}_{s_{1}}(k)\hat{\nu}^{\theta}_{s_{2}}(k^{\prime})]\Big{[}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}(e^{\frac{i\eta}{\varepsilon}[z_{1}g(\xi_{j},k)-z_{2}g(\zeta_{l},-k^{\prime})]}\psi(\xi_{j}-k,\zeta_{l}+k^{\prime})$
$\qquad\qquad+e^{-\frac{i\eta}{\varepsilon}[z_{1}g(\zeta_{l},-k)-z_{2}g(\xi_{j},k^{\prime})]}\psi(\xi_{j}-k^{\prime},\zeta_{l}+k))$
$-\sum\limits_{j\neq j^{\prime}}e^{\frac{i\eta}{\varepsilon}[z_{1}g(\xi_{j},k)+z_{2}g(\xi_{j^{\prime}},k^{\prime})]}\psi(\xi_{j}-k,\xi_{j^{\prime}}-k^{\prime})-\sum\limits_{l\neq l^{\prime}}e^{-\frac{i\eta}{\varepsilon}[z_{1}g(\zeta_{l},-k)+z_{2}g(\zeta_{l^{\prime}},-k^{\prime})]}\psi(\zeta_{l}+k,\zeta_{l^{\prime}}+k^{\prime})$
$-\sum\limits_{j=1}^{p}e^{\frac{i\eta}{\varepsilon}[z_{1}g(\xi_{j},k)+z_{2}g(\xi_{j}-k,k^{\prime})]}\psi(\xi_{j}-k-k^{\prime})-\sum\limits_{l=1}^{q}e^{-\frac{i\eta}{\varepsilon}[z_{1}g(\zeta_{l},-k)+z_{2}g(\zeta_{l}+k,-k^{\prime})]}\psi(\zeta_{l}+k+k^{\prime})\Big{]}\frac{\mathrm{d}k\mathrm{d}k^{\prime}}{(2\pi)^{2d}}\,.$
Using the stationarity property of $\nu$: $\mathbb{E}[\hat{\nu}^{\theta}_{s_{1}}(k)\hat{\nu}^{\theta}_{s_{2}}(k^{\prime})]=(2\pi)^{d}\hat{\mathfrak{C}}^{\theta}(s_{1}-s_{2},k)\delta(k+k^{\prime})$, where $\mathfrak{C}^{\theta}$ is defined in (12), we deduce from the above expression that
$\mathbb{E}[L^{\theta}(z_{1},\nu^{\theta}_{s_{1}})L^{\theta}(z_{2},\nu^{\theta}_{s_{2}})]\psi$ (44)
$=-\int\limits_{\mathbb{R}^{d}}\hat{\mathfrak{C}}^{\theta}(s_{1}-s_{2},k)\Big{[}\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}(e^{\frac{i\eta}{\varepsilon}[z_{1}g(\xi_{j},k)-z_{2}g(\zeta_{l},k)]}+e^{-\frac{i\eta}{\varepsilon}[z_{1}g(\zeta_{l},k)-z_{2}g(\xi_{j},k)]})\psi(\xi_{j}-k,\zeta_{l}-k)$
$-\sum\limits_{j\neq j^{\prime}}e^{\frac{i\eta}{\varepsilon}[z_{1}g(\xi_{j},k)+z_{2}g(\xi_{j^{\prime}},-k)]}\psi(\xi_{j}-k,\xi_{j^{\prime}}+k)-\sum\limits_{l\neq l^{\prime}}e^{-\frac{i\eta}{\varepsilon}[z_{1}g(\zeta_{l},k)+z_{2}g(\zeta_{l^{\prime}},-k)]}\psi(\zeta_{l}-k,\zeta_{l^{\prime}}+k)$
$-\sum\limits_{j=1}^{p}e^{\frac{i\eta}{\varepsilon}[z_{1}g(\xi_{j},k)+z_{2}g(\xi_{j}-k,-k)]}\psi-\sum\limits_{l=1}^{q}e^{-\frac{i\eta}{\varepsilon}[z_{1}g(\zeta_{l},k)+z_{2}g(\zeta_{l}-k,-k)]}\psi\Big{]}\frac{\mathrm{d}k}{(2\pi)^{d}}.$
From (41) and the above calculation, we thus have that
$\mathcal{L}^{\theta}(z)\psi$ $=\frac{1}{\eta\varepsilon\theta}\int\limits_{\mathbb{R}^{d}}\int\limits_{z}^{\infty}\hat{\mathfrak{C}}^{\theta}(z-s,k)\Big{[}2\sum\limits_{j=1}^{p}\sum\limits_{l=1}^{q}e^{\frac{i\eta z}{\varepsilon}[g(\xi_{j},k)-g(\zeta_{l},k)]}\psi(\xi_{j}-k,\zeta_{l}-k)$
$-\sum\limits_{j\neq j^{\prime}}e^{\frac{i\eta z}{\varepsilon}[g(\xi_{j},k)+g(\xi_{j^{\prime}},-k)]}\psi(\xi_{j}-k,\xi_{j^{\prime}}+k)-\sum\limits_{l\neq l^{\prime}}e^{-\frac{i\eta z}{\varepsilon}[g(\zeta_{l},k)+g(\zeta_{l^{\prime}},-k)]}\psi(\zeta_{l}-k,\zeta_{l^{\prime}}+k)$
$-\sum\limits_{j=1}^{p}e^{\frac{i\eta z}{\varepsilon}[g(\xi_{j},k)+g(\xi_{j}-k,-k)]}\psi-\sum\limits_{l=1}^{q}e^{-\frac{i\eta z}{\varepsilon}[g(\zeta_{l},k)+g(\zeta_{l}-k,-k)]}\psi\Big{]}\frac{\mathrm{d}s\mathrm{d}k}{(2\pi)^{d}}\,.$
This yields (38) after a change of variables $\frac{\eta(z-s)}{\varepsilon\theta}\to s$ and noting that
$g(\xi,k)+g(\xi-k,-k)=|\xi|^{2}-|\xi-k|^{2}+|\xi-k|^{2}-|\xi-k+k|^{2}=0\,.$
The bound (39) and the proof of the theorem require lengthy estimates obtained in Lemmas 4.7, 4.8 and 4.9 below. ∎
We note that $\mathcal{L}^{\theta}(z)$ is precisely the integral operator of the phase compensated moments described in [5] under the Itô-Schrödinger white noise regime. The second term in the remainder comes from a mismatch in phases evaluated at different times. The third term contains correlations between points in $z$ located far apart and is small by our assumptions of decorrelations in $z$. Finding bounds for the first term is non-trivial and we will use the Duhamel representation of the solution here. We will show that this term involves pairings not consistent with the time ordering in the simplex $0\leq s^{\prime}_{1}\leq\cdots\leq s^{\prime}_{2N}\leq s^{\prime}\leq s\leq z$ (and similar to the non ladder diagrams in [4]).
Before proving (39), we conclude the proofs of Theorems 3.1 and 3.2. We first note that the above result translates in the physical domain for $\mu^{\theta}_{p,q}$ defined in (7) as follows.
Remark 4.4.
The $p+q$th moment of $u^{\theta}$ can be decomposed as
$\mu_{p,q}^{\theta}(z,X,Y)=\tilde{\mathcal{B}}^{\theta}_{p,q}(z,X,Y)+\tilde{\mathscr{R}}_{p,q}^{\theta}(z,X,Y)\,,$
with ~$\tilde{\mathcal{B}}^{\theta}_{p,q}$ given by
~$\tilde{\mathcal{B}}^{\theta}_{p,q}(z,X,Y)=\int\limits_{\mathbb{R}^{(p+q)d}}\Pi^{\theta}(z,v)\mathcal{B}^{\theta}_{p,q}(z,v)e^{i(\sum_{j=1}^{p}\xi_{j}\cdot x_{j}-\sum_{l=1}^{q}\zeta_{l}\cdot y_{l})}\prod\limits_{j=1}^{p}\frac{\mathrm{d}\xi_{j}}{(2\pi)^{d}}\prod\limits_{l=1}^{q}\frac{\mathrm{d}\zeta_{l}}{(2\pi)^{d}}\,,$
where $\mathcal{B}^{\theta}_{p,q}$ is given by (37) and ~$\tilde{\mathscr{R}}_{p,q}^{\theta}$ satisfies
~$\|\tilde{\mathscr{R}}_{p,q}^{\theta}(z,X,Y)\|_{\infty}\leq c(p,q,z)\frac{\theta}{\eta^{2}}e^{\frac{\lambda[R,p,q,z,0]}{\eta^{2}}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{2}\psi^{\theta}(0,v)\|\,.$
Convergence of moments: From paraxial to complex Gaussian.
Let $\gamma=(j,l)$ denote a pairing with $1\leq j\leq p,1\leq l\leq q$. We denote by $\Lambda_{\kappa}$, the set of all pairings $(j,l)$, $1\leq j\leq p,1\leq l\leq q$ such that if $\gamma,\gamma^{\prime}\in\Lambda_{\kappa}$, then $\gamma_{1,2}\neq\gamma^{\prime}_{1,2}$. For a given arrangement $\kappa$, let $m(\kappa)$ denote the number of such pairings in the set $\Lambda_{\kappa}$ and $K=\sum m(\kappa)=\sum\limits_{m=1}^{p\wedge q}\binom{p}{m}\binom{q}{m}m!$ denote the total number of all such distinct pairings.
Define $U^{\theta}_{\gamma}(z)$ to be a bounded operator on $\mathcal{M}_{B}(\mathbb{R}^{(p+q)d})$ such that $\rho(z,v)=[U^{\theta}_{\gamma}(z)\rho(0)](v),\rho(0,v)\in\mathcal{M}_{B}(\mathbb{R}^{(p+q)d})$ satisfies
$\partial_{z}\rho=\frac{1}{\eta^{2}}\int\limits_{\mathbb{R}^{d}}\hat{R}(k)e^{\frac{i\eta z}{\varepsilon}[g(\xi_{j},k)-g(\zeta_{l},k)]}\rho(z,\xi_{j}-k,\zeta_{l}-k)\frac{\mathrm{d}k}{(2\pi)^{d}}\,.$
Also, let $U_{\eta}(z)=e^{-\frac{R(0)z}{\eta^{2}}}$.
Theorem 4.5.
The $p+q$th moments of the phase compensated field satisfy the decomposition
$\Psi^{\theta}_{p,q}(z,v)=[N_{p,q}^{\theta}(z)\psi^{\theta}(0)](v)+E^{\theta}_{p,q}(z,v)\,,$ (45)
where $N_{p,q}^{\theta}(z)$ is a bounded operator on $\mathcal{M}_{B}(\mathbb{R}^{(p+q)d})$ given by
$N_{p,q}^{\theta}(z)=U_{\eta}^{\frac{p+q}{2}}(z)\sum\limits_{\kappa=1}^{K}\prod\limits_{\gamma\in\Lambda_{\kappa}}(U^{\theta}_{\gamma}(z)-\mathbb{I})+U_{\eta}^{\frac{p+q}{2}}(z)\,,$ (46)
and $E^{\theta}_{p,q}$ satisfies
$\|E_{p,q}^{\theta}(z,v)\|\leq c(p,q,z)(\varepsilon^{\frac{1}{3}}+\theta^{\frac{1}{2}})\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{2}\psi^{\theta}(0)\|\,.$ (47)
Proof.
From Theorem 4.3, we have that
$\Psi_{p,q}^{\theta}(z,v)=\mathcal{B}^{\theta}_{p,q}(z,v)+\mathscr{R}^{\theta}_{p,q}(z,v)\,,$
with $\mathcal{B}^{\theta}_{p,q}$ as in (37) and $\mathscr{R}^{\theta}$ satisfying the bound (39). Note that $\mathcal{B}_{p,q}^{\theta}$ follows the same evolution equation as that of the phase compensated moments of the Itô-Schrödinger equation. Due to this, the results in [5] apply and using [5, Theorem 4.2], we have that
$\mathcal{B}^{\theta}_{p,q}(z,v)=[N^{\theta}_{p,q}(z)\psi^{\theta}(0)](v)+\mathcal{E}_{p,q}^{\theta}(z,v)\,,$
with $N_{p,q}^{\theta}(z)$ as defined in (46) and $\mathcal{E}_{p,q}^{\theta}$ satisfying
$\|\mathcal{E}_{p,q}^{\theta}(z,v)\|\leq c(p,q,z)\varepsilon^{\frac{1}{3}}\|\prod\limits_{m=1}^{p+q}\psi^{\theta}(0)\|\,.$
This gives
$E^{\theta}_{p,q}(z,v)=\mathscr{R}_{p,q}^{\theta}(z,v)+\mathcal{E}_{p,q}^{\theta}(z,v)$
with $E^{\theta}_{p,q}$ satisfying the bound in (47). ∎
Define the inverse (phase re-compensated) Fourier transform
$\tilde{N}^{\theta}_{p,q}(z,X,Y)=\int\limits_{\mathbb{R}^{(p+q)d}}\Pi^{\theta}(z,v)[N_{p,q}^{\theta}(z)\psi^{\theta}(0)](v)e^{i(\sum_{j=1}^{p}\xi_{j}\cdot x_{j}-\sum_{l=1}^{q}\zeta_{l}\cdot y_{l})}\prod\limits_{j=1}^{p}\frac{\mathrm{d}\xi_{j}}{(2\pi)^{d}}\prod\limits_{l=1}^{q}\frac{\mathrm{d}\zeta_{l}}{(2\pi)^{d}}\,,$ (48)
with $\tilde{E}^{\theta}_{p,q}(z,X,Y)$ defined in a similar manner for $E^{\theta}_{p,q}$. Then we have the following direct consequence of the above results in physical domain:
Theorem 4.6.
The $p+q$th moments of $u^{\theta}$ can be written as
$\mu^{\theta}_{p,q}(z,X,Y)=\tilde{N}_{p,q}^{\theta}(z,X,Y)+\tilde{E}^{\theta}_{p,q}(z,X,Y)\,,$
with $\tilde{N}_{p,q}^{\theta}$ as defined in (48) and
$\|\tilde{E}_{p,q}^{\theta}(z,X,Y)\|_{\infty}\leq c(p,q,z)(\varepsilon^{\frac{1}{3}}+\theta^{\frac{1}{2}})\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{2}\psi^{\theta}(0)\|\,.$
Proof of Theorems 3.1 and 3.2.
These are direct consequences of the estimate in Theorem 4.6 combined with the corresponding results in the Itô-Schrödinger regime [5, Theorems 2.2 & 2.4]. ∎
[…]
Control of the error terms in (39).
We now state and prove the lemmas allowing us to conclude the proof of Theorem 4.3.
Lemma 4.7.
We have the first estimate:
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{1}(z,v)\|\leq c_{1}(p,q,z)\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\,,$ (49)
where $\lambda$ is defined in (36), $0\leq\alpha\leq\frac{d_{0}+1}{p+q}$ and $c_{1}$ is a constant independent of $\theta$.
Proof.
Plugging the chaos expansion (29)-(31) in (43) and using the Gaussian property
$\mathbb{E}[\hat{\nu}^{\theta}_{s}(k)\hat{\nu}^{\theta}_{s^{\prime}}(k^{\prime})\prod\limits_{j=1}^{2N}\hat{\nu}^{\theta}_{s_{j}}(k_{j})]-\mathbb{E}[\hat{\nu}^{\theta}_{s}(k)\hat{\nu}^{\theta}_{s^{\prime}}(k^{\prime})]\mathbb{E}[\prod\limits_{j=1}^{2N}\hat{\nu}^{\theta}_{s_{j}}(k_{j})]$
$=\sum\limits_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}\mathbb{E}[\hat{\nu}^{\theta}_{s}(k)\hat{\nu}^{\theta}_{s_{m}}(k_{m})]\mathbb{E}[\hat{\nu}^{\theta}_{s^{\prime}}(k^{\prime})\hat{\nu}^{\theta}_{s_{n}}(k_{n})]\mathbb{E}[\prod_{\begin{subarray}{c}j=1\\ j\neq m,n\end{subarray}}^{2N}\hat{\nu}^{\theta}_{s_{j}}(k_{j})]\,,$
we deduce that
$\mathscr{R}^{\theta}_{1}(z,v)$ $=\sum\limits_{N\geq 1}\frac{(-1)^{N+1}}{(\eta\varepsilon\theta)^{N+1}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2N}_{<}}\int\limits_{\mathbb{R}^{(2N+2)d}}\sum_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}\mathbb{E}[\hat{\nu}^{\theta}_{s}(k)\hat{\nu}^{\theta}_{s_{m}}(k_{m})]\mathbb{E}[\hat{\nu}^{\theta}_{s^{\prime}}(k^{\prime})\hat{\nu}^{\theta}_{s_{n}}(k_{n})]$
×$\times\mathbb{E}[\prod_{\begin{subarray}{c}j=1\\ j\neq m,n\end{subarray}}^{2N}\hat{\nu}^{\theta}_{s_{j}}(k_{j})]\sum\limits_{j=1}^{(p+q)^{2N+2}}e^{\frac{i\eta}{\varepsilon}G_{j}(\vec{s},\vec{k},v)}\psi^{\theta}(0,v-A_{j}\vec{k})\prod\limits_{j=1}^{2N+2}\frac{\mathrm{d}k_{j}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\mathrm{d}s^{\prime}\mathrm{d}s\,.$
Again using the Gaussian structure of the potential, we deduce using (30) that
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{1}(z,v)\|$ $\leq\sum\limits_{N\geq 1}\frac{1}{(\eta\varepsilon\theta)^{N+1}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2N}_{<}}\int\limits_{\mathbb{R}^{(N+1)d}}\sum_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}\sum\limits_{P^{2N-2}}\sum\limits_{j=1}^{(p+q)^{2N+2}}\|\prod\limits_{m^{\prime}=1}^{p+q}\langle v_{m^{\prime}}\rangle^{\alpha}\psi^{\theta}(0,v-A_{j}^{\prime}\vec{k})\|$
×^$\times\hat{\mathfrak{C}}^{\theta}(s-s_{m},k)\hat{\mathfrak{C}}^{\theta}(s^{\prime}-s_{n},k^{\prime})\frac{\mathrm{d}k\mathrm{d}k^{\prime}}{(2\pi)^{2d}}\prod\limits_{\begin{subarray}{c}(j^{\prime},l^{\prime})\in P^{2N-2}\\ j^{\prime},l^{\prime}\neq m,n\end{subarray}}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\mathrm{d}s^{\prime}\mathrm{d}s\,.$
Note that as in (34)
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v-A_{j}^{\prime}\vec{k})\|$ $=\|\prod\limits_{m=1}^{p+q}\langle v_{m}+\sum\limits_{l=1}^{N+1}[A_{j}^{\prime}]_{m,l}k_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|\,,$
with $k_{N},k_{N+1}$ taken as $k^{\prime},k$ respectively. From Lemma 4.1, $|[A_{j}^{\prime}]_{m,l}|\leq C$ and we have as in (35) that
$1\leq\langle v_{m}+\sum\limits_{l=1}^{N+1}[A_{j}^{\prime}]_{m,l}k_{m}\rangle\leq C^{N+1}\langle v_{m}\rangle\prod\limits_{l=1}^{N+1}\langle k_{l}\rangle.$
Using this, we have that $\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{1}(z,v)\|$ is bounded by
$\sum\limits_{N\geq 1}\frac{C^{(N+1)(p+q)}(p+q)^{2N+2}}{(\eta\varepsilon\theta)^{N+1}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|$
×$\times\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{\mathbb{R}^{2d}}\sum_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}(\langle k\rangle\langle k^{\prime}\rangle)^{(p+q)\alpha}\hat{\mathfrak{C}}^{\theta}(s-s_{m},k)\hat{\mathfrak{C}}^{\theta}(s^{\prime}-s_{n},k^{\prime})\frac{\mathrm{d}k\mathrm{d}k^{\prime}}{(2\pi)^{2d}}\mathrm{d}s^{\prime}\mathrm{d}s$
×$\times\int\limits_{[0,s^{\prime}]^{2N}_{<}}\int\limits_{\mathbb{R}^{(N-1)d}}\sum\limits_{P^{2N-2}}\prod\limits_{\begin{subarray}{c}(j^{\prime},l^{\prime})\in P^{2N-2}\\ j^{\prime},l^{\prime}\neq m,n\end{subarray}}\langle k_{j^{\prime}}\rangle^{(p+q)\alpha}\hat{\mathfrak{C}}^{\theta}(s_{j^{\prime}}-s_{l^{\prime}},k_{j^{\prime}})\frac{\mathrm{d}k_{j^{\prime}}}{(2\pi)^{d}}\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\,.$
Let $\mathfrak{C}_{n,\alpha}(s)=\int\limits_{\mathbb{R}^{d}}\langle k\rangle^{n\alpha}\hat{\mathfrak{C}}(s,k)\frac{\mathrm{d}k}{(2\pi)^{d}}$ with $\mathfrak{C}^{\theta}_{n,\alpha}(s)=\mathfrak{C}_{n,\alpha}\big{(}\frac{\eta s}{\varepsilon\theta}\big{)}$. Then we have that
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{1}(z,v)\|\ \leq\ \|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|\sum\limits_{N\geq 1}\frac{C^{(N+1)(p+q)}(p+q)^{2N+2}}{(\eta\varepsilon\theta)^{N+1}}$
×$\times\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2N}_{<}}\sum_{\begin{subarray}{c}m,n=1\\ n\neq m\end{subarray}}^{2N}\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{m})\mathfrak{C}_{p+q,\alpha}^{\theta}(s^{\prime}-s_{n})\sum\limits_{P^{2N-2}}\prod\limits_{\begin{subarray}{c}(j,l)\in P^{2N-2}\\ j,l\neq m,n\end{subarray}}\mathfrak{C}_{p+q,\alpha}^{\theta}(s_{j}-s_{l})\prod\limits_{j=1}^{2N}\mathrm{d}s_{j}\mathrm{d}s^{\prime}\mathrm{d}s\,.$
Again, symmetrizing the simplex leads to an expression of the form
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{1}(z,v)\|\leq\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|\sum\limits_{N\geq 1}\frac{C^{(N+1)(p+q)}(p+q)^{2N+2}}{(N-1)!}$ (50)
×$\times\Big{(}\frac{1}{\eta\varepsilon\theta}\int\limits_{[0,z]^{2}}\mathfrak{C}^{\theta}_{p+q,\alpha}(s-s^{\prime})\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}^{N-1}\Big{(}\frac{1}{(\eta\varepsilon\theta)^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2}}\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{1})\mathfrak{C}_{p+q,\alpha}^{\theta}(s^{\prime}-s_{2})\mathrm{d}s_{2}\mathrm{d}s_{1}\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}.$
As before, we have that
$\frac{1}{\eta\varepsilon\theta}\int\limits_{[0,z]^{2}}\mathfrak{C}^{\theta}_{p+q,\alpha}(s-s^{\prime})\mathrm{d}s^{\prime}\mathrm{d}s\leq\frac{z}{\eta^{2}}\int\limits_{\mathbb{R}}\mathfrak{C}_{p+q,\alpha}(s)\mathrm{d}s=\frac{z}{\eta^{2}}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|.$
On the other hand,
$\frac{1}{(\eta\varepsilon\theta)^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{[0,s^{\prime}]^{2}}\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{1})\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{2})\mathrm{d}s_{2}\mathrm{d}s_{1}\mathrm{d}s^{\prime}\mathrm{d}s$
$\leq\Big{(}\frac{1}{\eta\varepsilon\theta}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{0}^{s^{\prime}}\mathfrak{C}_{p+q,\alpha}^{\theta}(s-s_{1})\mathrm{d}s_{1}\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}\Big{(}\frac{1}{\eta^{2}}\int\limits_{\mathbb{R}}\mathfrak{C}_{p+q,\alpha}(s)\mathrm{d}s\Big{)}$
$=\Big{(}\frac{1}{\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{\frac{\eta(s-s^{\prime})}{\varepsilon\theta}}^{\frac{\eta s}{\varepsilon\theta}}\mathfrak{C}_{p+q,\alpha}(s_{1})\mathrm{d}s_{1}\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}\Big{(}\frac{1}{\eta^{2}}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|\Big{)}.$
From assumption (9), $\mathfrak{C}_{p+q,\alpha}(s)\leq\lambda_{1}(d_{0})e^{-\tau_{R}|s|}$. So, we deduce that the above integral is bounded as
$\Big{(}\frac{1}{\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{\frac{\eta(s-s^{\prime})}{\varepsilon\theta}}^{\frac{\eta s}{\varepsilon\theta}}\mathfrak{C}_{p+q,\alpha}(s_{1})\mathrm{d}s_{1}\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}\Big{(}\frac{1}{\eta^{2}}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|\Big{)}$
$\leq\Big{(}\frac{\lambda_{1}(d_{0})}{\tau_{R}\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{s}e^{\frac{\tau_{R}\eta}{\varepsilon\theta}(s-s^{\prime})}\mathrm{d}s^{\prime}\mathrm{d}s\Big{)}\Big{(}\frac{1}{\eta^{2}}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|\Big{)}\leq\lambda_{1}(d)\frac{\varepsilon\theta}{\tau_{R}^{2}\eta^{3}}\Big{(}\frac{z}{\eta^{2}}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|\Big{)}\,.$
Together, these lead to a bound of the form
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{1}(z,v)\|$ $\leq\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|\sum\limits_{N\geq 1}\frac{(p+q)^{2N+2}}{(N-1)!}C^{(N+1)(p+q)}\frac{z^{N}}{\eta^{2N}}\Big{(}\|\langle k\rangle^{(p+q)\alpha}\hat{R}(k)\|\Big{)}^{N}$
$\leq c_{1}(p,q,z)\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\,.$
This leads to (49) and concludes the derivation of the first estimate. ∎
Lemma 4.8.
We have the following second estimate:
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{2}(z,v)\|\leq c_{2}(p,q,z)\frac{\theta}{\eta^{2}}e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{1+\alpha}\psi^{\theta}(0)\|\,,$ (51)
with $\lambda$ as in (36) and $0\leq\alpha\leq\frac{d_{0}+1}{p+q}$. Here, $c_{2}$ is a constant independent of $\theta$.
Proof.
From (43) and (44), we have
$\mathscr{R}^{\theta}_{2}(z,v)=\frac{1}{\eta\varepsilon\theta}\int\limits_{0}^{z}\int\limits_{0}^{s}\int\limits_{\mathbb{R}^{d}}\hat{\mathfrak{C}}^{\theta}(s-s^{\prime},k)$
×$\times\Big{[}\sum\limits_{j,l}\big{(}2e^{\frac{i\eta}{\varepsilon}s^{\prime}[g(\xi_{j},k)-g(\zeta_{l},k)]}-e^{\frac{i\eta}{\varepsilon}[sg(\xi_{j},k)-s^{\prime}g(\zeta_{l},k)]}-e^{-\frac{i\eta}{\varepsilon}[sg(\zeta_{l},k)-s^{\prime}g(\xi_{j},k)]}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)$
$-\sum\limits_{j\neq j^{\prime}}\big{(}e^{\frac{i\eta s^{\prime}}{\varepsilon}[g(\xi_{j},k)+g(\xi_{j^{\prime}},-k)]}-e^{\frac{i\eta}{\varepsilon}[sg(\xi_{j},k)+s^{\prime}g(\xi_{j^{\prime}},-k)]}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\xi_{j^{\prime}}+k)$
$-\sum\limits_{l\neq l^{\prime}}\big{(}e^{-\frac{i\eta s^{\prime}}{\varepsilon}[g(\zeta_{l},k)+g(\zeta_{l^{\prime}},-k)]}-e^{-\frac{i\eta}{\varepsilon}[sg(\zeta_{l},k)+s^{\prime}g(\zeta_{l^{\prime}},-k)]}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime},\zeta_{l}-k,\zeta_{l^{\prime}}+k)$
$-\sum\limits_{j}\big{(}1-e^{\frac{i\eta}{\varepsilon}[sg(\xi_{j},k)+s^{\prime}g(\xi_{j}-k,-k)]}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime})-\sum\limits_{l}\big{(}1-e^{-\frac{i\eta}{\varepsilon}[sg(\zeta_{l},k)+s^{\prime}g(\zeta_{l}-k,-k)]}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime})\Big{]}\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s^{\prime}\mathrm{d}s\,.$
Switching the order of integration of $s^{\prime}$ and $s$ and performing a change of variables $\frac{\eta(s-s^{\prime})}{\varepsilon\theta}\to s$ gives
$\mathscr{R}^{\theta}_{2}(z,v)$ $=\frac{1}{\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{\frac{\eta(z-s^{\prime})}{\varepsilon\theta}}\int\limits_{\mathbb{R}^{d}}\hat{\mathfrak{C}}(s,k)$
×$\times\Big{[}\sum\limits_{j,l}e^{\frac{i\eta}{\varepsilon}s^{\prime}[g(\xi_{j},k)-g(\zeta_{l},k)]}\big{(}2-e^{i\theta sg(\xi_{j},k)}-e^{-i\theta sg(\zeta_{l},k)}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)$
$-\sum\limits_{j\neq j^{\prime}}e^{\frac{i\eta s^{\prime}}{\varepsilon}[g(\xi_{j},k)+g(\xi_{j^{\prime}},-k)]}\big{(}1-e^{i\theta sg(\xi_{j},k)}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\xi_{j^{\prime}}+k)$
$-\sum\limits_{l\neq l^{\prime}}e^{-\frac{i\eta s^{\prime}}{\varepsilon}[g(\zeta_{l},k)+g(\zeta_{l^{\prime}},-k)]}\big{(}1-e^{-i\theta sg(\zeta_{l},k)}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime},\zeta_{l}-k,\zeta_{l^{\prime}}+k)$
$-\sum\limits_{j}\big{(}1-e^{i\theta sg(\xi_{j},k)}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime})-\sum\limits_{l}\big{(}1-e^{-i\theta sg(\zeta_{l},k)}\big{)}\mathbb{E}\psi^{\theta}(s^{\prime})\Big{]}\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s\mathrm{d}s^{\prime}\,.$
From the above expression, we have that
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{2}(z,v)\|$
$\leq\frac{\theta}{\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{\frac{\eta(z-s^{\prime})}{\varepsilon\theta}}\int\limits_{\mathbb{R}^{d}}|s|\hat{\mathfrak{C}}(s,k)\int\limits_{\mathbb{R}^{(p+q)d}}\Big{[}\sum\limits_{j,l}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\big{(}|g(\xi_{j},k)|+|g(\zeta_{l},k)|\big{)}|\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)|$
$+\sum\limits_{j\neq j^{\prime}}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}|g(\xi_{j},k)||\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\xi_{j^{\prime}}+k)|+\sum\limits_{l\neq l^{\prime}}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}|g(\zeta_{l},k)||\mathbb{E}\psi^{\theta}(s^{\prime},\zeta_{l}-k,\zeta_{l^{\prime}}+k)|$
$+\sum\limits_{j}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}|g(\xi_{j},k)||\mathbb{E}\psi^{\theta}(s^{\prime})|+\sum\limits_{l}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}|g(\zeta_{l},k)||\mathbb{E}\psi^{\theta}(s^{\prime})|\Big{]}\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s\mathrm{d}s^{\prime}\,.$
For a fixed $(j,l)$ in the first sum, we make a change of variables $\xi_{j}-k\to\xi_{j},\zeta_{l}-k\to\zeta_{l}$. This transformation leads to a product of the form
$\Big{(}\prod_{\begin{subarray}{c}m=1\\ m\neq j,l\end{subarray}}^{2N}\langle v_{m}\rangle^{\alpha}\Big{)}\langle\xi_{j}+k\rangle^{\alpha}\langle\zeta_{l}+k\rangle^{\alpha}\Big{(}|g(\xi_{j}+k,k)|+|g(\zeta_{l}+k,k)|\Big{)}.$
Since $1\leq\langle\xi_{j}+k\rangle\langle\zeta_{l}+k\rangle\leq 2\langle\xi_{j}\rangle\langle\zeta_{l}\rangle\langle k\rangle^{2}$ and
$|g(\xi_{j}+k,k)|+|g(\zeta_{l}+k,k)|$ $\leq 2|k|^{2}+2|k|(|\xi_{j}|+|\zeta_{l}|)\leq 2\langle k\rangle^{2}+2\langle k\rangle(\langle\xi_{j}\rangle+\langle\zeta_{l}\rangle)\leq C\langle k\rangle^{2}\langle\xi_{j}\rangle\langle\zeta_{l}\rangle\,,$
we have that
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{2}(z,v)\|$
$\leq C\frac{\theta}{\eta^{2}}\int\limits_{0}^{z}\int\limits_{0}^{\frac{\eta(z-s^{\prime})}{\varepsilon\theta}}\int\limits_{\mathbb{R}^{d}}|s|\langle k\rangle^{2+2\alpha}\hat{\mathfrak{C}}(s,k)\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\langle v_{n}\rangle\langle v_{n^{\prime}}\rangle\Big{(}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\Big{)}\mathbb{E}\psi^{\theta}(s^{\prime},v)\|\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s\mathrm{d}s^{\prime}$
$\leq Cz\frac{\theta}{\eta^{2}}\sup\limits_{0\leq s\leq z}\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\langle v_{n}\rangle\langle v_{n^{\prime}}\rangle\Big{(}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\Big{)}\mathbb{E}\psi^{\theta}(s,v)\|\int\limits_{0}^{\infty}\int\limits_{\mathbb{R}^{d}}|s|\langle k\rangle^{2+2\alpha}\hat{\mathfrak{C}}(s,k)\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s\,.$
The bound in (51) now follows from assumption (9) and using Lemma 4.2 applied to
$\|\sum\limits_{\begin{subarray}{c}n,n^{\prime}=1\\ n\neq n^{\prime}\end{subarray}}^{p+q}\langle v_{n}\rangle\langle v_{n^{\prime}}\rangle\Big{(}\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\Big{)}\mathbb{E}\psi^{\theta}(s,v)\|.$
This concludes the proof of the Lemma. ∎
The final error term in (39) is bounded as follows:
Lemma 4.9.
We have the following third estimate:
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|\leq c_{3}(p,q)\frac{\varepsilon\theta}{\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\psi^{\theta}(0,v)\|e^{\frac{\lambda[R,p,q,z,\alpha]}{\eta^{2}}}\,,$ (52)
with $\lambda$ as in (36) and $0\leq\alpha\leq\frac{d_{0}+1}{p+q}$.
Proof.
From (43) and (44), we have that
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|$ $\leq\frac{1}{\eta\varepsilon\theta}\int\limits_{z}^{\infty}\int\limits_{0}^{z}\int\limits_{\mathbb{R}^{d}}\hat{\mathfrak{C}}^{\theta}(s-s^{\prime},k)\Big{[}2\sum\limits_{j,l}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)\|$
$+\sum\limits_{j\neq j^{\prime}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\xi_{j^{\prime}}+k)\|+\sum\limits_{l\neq l^{\prime}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s^{\prime},\zeta_{l}-k,\zeta_{l^{\prime}}+k)\|$
$+(p+q)\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s^{\prime},v)\|\Big{]}\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s^{\prime}\mathrm{d}s\,.$
As before, making an appropriate change of variables gives for instance
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s^{\prime},\xi_{j}-k,\zeta_{l}-k)\|=\|\prod\limits_{m=1,m\neq j,p+l}^{p+q}\langle v_{m}\rangle^{\alpha}\langle\xi_{j}+k\rangle^{\alpha}\langle\zeta_{l}+k\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s^{\prime},v)\|,$
which is bounded by $C\langle k\rangle^{2\alpha}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s^{\prime},v)\|$. This gives
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|$ $\leq\frac{(p+q)^{2}}{\eta\varepsilon\theta}\sup\limits_{0\leq s\leq z}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s,v)\|\int\limits_{z}^{\infty}\int\limits_{0}^{z}\int\limits_{\mathbb{R}^{d}}\langle k\rangle^{2\alpha}\hat{\mathfrak{C}}^{\theta}(s-s^{\prime},k)\frac{\mathrm{d}k}{(2\pi)^{d}}\mathrm{d}s^{\prime}\mathrm{d}s\,.$
Since $\int\limits_{\mathbb{R}^{d}}\langle k\rangle^{2\alpha}\hat{\mathfrak{C}}(s,k)\frac{\mathrm{d}k}{(2\pi)^{d}}\leq Ce^{-\tau_{R}|s|}$, after a change of variables $\frac{\eta(s-s^{\prime})}{\varepsilon\theta}\to s^{\prime}$, we have
$\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathscr{R}^{\theta}_{3}(z,v)\|$ $\leq C\frac{(p+q)^{2}}{\eta^{2}}\sup\limits_{0\leq s\leq z}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s,v)\|\int\limits_{z}^{\infty}\int\limits_{\frac{\eta(s-z)}{\varepsilon\theta}}^{\frac{\eta s}{\varepsilon\theta}}e^{-\tau_{R}|s|}\mathrm{d}s^{\prime}\mathrm{d}s$
$\leq C\frac{(p+q)^{2}}{\tau_{R}\eta^{2}}\sup\limits_{0\leq s\leq z}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s,v)\|\int\limits_{z}^{\infty}e^{-\frac{\tau_{R}\eta s}{\varepsilon\theta}}\mathrm{d}s$
$\leq C(p+q)^{2}\frac{\varepsilon\theta}{\tau_{R}^{2}\eta^{3}}\|\prod\limits_{m=1}^{p+q}\langle v_{m}\rangle^{\alpha}\mathbb{E}\psi^{\theta}(s,v)\|\,.$
The bound in (52) now follows from using Lemma 4.2 in the above expression. This concludes the derivation of the third estimate and hence the proof of Theorem 4.3. ∎
4.3 Proof of the tightness result
Proof of Theorem 3.3.
We follow a similar strategy to the proof of [5, Theorem 2.7]. Let $\delta_{h}\phi^{\theta}(z,r,x):=\phi^{\theta}(z,r,x+h)-\phi^{\theta}(z,r,x)$. Then,
$\mathbb{E}|\phi^{\theta}(z,r,x+h)-\phi^{\theta}(z,r,x)|^{2n}=\mathbb{E}|\delta_{h}\phi^{\theta}(z,r,x)|^{2n}$
$=\int\limits_{\mathbb{R}^{2nd}}\prod\limits_{j=1}^{n}[(e^{i\xi_{j}\cdot h}-1)(e^{-i\zeta_{l}\cdot h}-1)]\mathbb{E}[\prod\limits_{j=1}^{n}\hat{\phi}^{\theta}(z,r,\xi_{j})\prod\limits_{l=1}^{n}\hat{\phi}^{\theta*}(z,r,\zeta_{l})]e^{i\sum_{j=1}^{n}(\xi_{j}-\zeta_{j})\cdot x}\prod\limits_{j=1}^{n}\frac{\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}}{(2\pi)^{2d}}\,.$
Note that
$\hat{\phi}^{\theta}(z,r,\xi)=\int\limits_{\mathbb{R}^{d}}u^{\theta}(z,\varepsilon^{-\beta}r+\eta x)e^{-i\xi\cdot x}\mathrm{d}x=\frac{1}{\eta^{d}}\hat{u}^{\theta}(z,\eta^{-1}\xi)e^{\frac{i}{\varepsilon^{\beta}\eta}r\cdot\xi}\,.$
From Theorem 4.3, we deduce that
$\mathbb{E}[\prod\limits_{j=1}^{n}\hat{\phi}^{\theta}(z,r,\xi_{j})\prod\limits_{l=1}^{n}\hat{\phi}^{\theta*}(z,r,\zeta_{l})]$ $=\frac{1}{\eta^{2nd}}e^{\frac{i}{\varepsilon^{\beta}\eta}r\cdot\sum_{j=1}^{n}(\xi_{j}-\zeta_{j})}\hat{\mu}^{\theta}_{n,n}(z,\eta^{-1}v)$ (53)
$=\frac{1}{\eta^{2nd}}e^{\frac{i}{\varepsilon^{\beta}\eta}r\cdot\sum_{j=1}^{n}(\xi_{j}-\zeta_{j})}\Pi^{\theta}_{n,n}(z,\eta^{-1}v)[\mathcal{B}^{\theta}_{n,n}(z,\eta^{-1}v)+\mathscr{R}^{\theta}_{n,n}(z,\eta^{-1}v)]\,.$
Note that
$\Pi^{\theta}_{n,n}(z,v)\mathcal{B}^{\theta}_{n,n}(z,v)=\mathbb{E}[\prod\limits_{j=1}^{n}\hat{u}^{\varepsilon}(z,\xi_{j})\hat{u}^{\varepsilon*}(z,\zeta_{j})]\,,$
where $\hat{u}^{\varepsilon}$ is the Fourier transform of $u^{\varepsilon}$, the solution to the Itô-Schrödinger equation (5) recast when $2k_{0}=1,R\to 16R$ as
$\mathrm{d}u^{\varepsilon}=\frac{i\eta}{\varepsilon}\Delta_{x}u^{\varepsilon}\mathrm{d}z-\frac{R(0)}{2\eta^{2}}u^{\varepsilon}\mathrm{d}z+\frac{i}{\eta}u^{\varepsilon}\mathrm{d}B,\quad u^{\varepsilon}(0,x)=u_{0}^{\varepsilon}(x)\,.$
So the first term in the sum (53) corresponds to the moments of the Itô-Schrödinger equation. Performing an inverse Fourier transform on this gives, using the shorthand notation $\mathrm{d}v=\prod_{j=1}^{n}(2\pi)^{-2d}\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}$:
$\frac{1}{\eta^{2nd}}\int\limits_{\mathbb{R}^{2nd}}\prod\limits_{j=1}^{n}[(e^{i\xi_{j}\cdot h}-1)(e^{-i\zeta_{l}\cdot h}-1)]e^{i\sum_{j=1}^{n}(\xi_{j}-\zeta_{j})\cdot x}e^{\frac{i}{\varepsilon^{\beta}\eta}r\cdot\sum_{j=1}^{n}(\xi_{j}-\zeta_{j})}\mathbb{E}[\prod\limits_{j=1}^{n}\hat{u}^{\varepsilon}(z,\eta^{-1}\xi_{j})\hat{u}^{\varepsilon*}(z,\eta^{-1}\zeta_{j})]\mathrm{d}v$
$=\int\limits_{\mathbb{R}^{2nd}}\prod\limits_{j=1}^{n}[(e^{i\xi_{j}\cdot(\frac{r}{\varepsilon^{\beta}}+\eta x+\eta h)}-e^{i\xi_{j}\cdot(\frac{r}{\varepsilon^{\beta}}+\eta x)})(e^{-i\zeta_{l}\cdot(\frac{r}{\varepsilon^{\beta}}+\eta x+\eta h)}-e^{-i\zeta_{l}\cdot(\frac{r}{\varepsilon^{\beta}}+\eta x)})]\mathbb{E}[\prod\limits_{j=1}^{n}\hat{u}^{\varepsilon}(z,\xi_{j})\hat{u}^{\varepsilon*}(z,\zeta_{j})]\mathrm{d}v$
$=\mathbb{E}|\phi^{\varepsilon}(z,r,x+h)-\phi^{\varepsilon}(z,r,x)|^{2n}:=\mathbb{E}|\delta_{h}\phi^{\varepsilon}(z,r,x)|^{2n}\,,$
where $\phi^{\varepsilon}(z,r,x)=u^{\varepsilon}(z,\varepsilon^{-\beta}r+\eta x)$ as in (14). (53) can now be written as
$\mathbb{E}|\delta_{h}\phi^{\theta}(z,r,x)|^{2n}=\mathbb{E}|\delta_{h}\phi^{\varepsilon}(z,r,x)|^{2n}+$
$\frac{1}{\eta^{2nd}}\int\limits_{\mathbb{R}^{2nd}}\prod\limits_{j=1}^{n}[(e^{i\xi_{j}\cdot h}-1)(e^{-i\zeta_{l}\cdot h}-1)]e^{\frac{i}{\varepsilon^{\beta}\eta}r\cdot\sum_{j=1}^{n}(\xi_{j}-\zeta_{j})}\Pi^{\theta}_{n,n}(z,\eta^{-1}v)\mathscr{R}^{\theta}_{n,n}(z,\eta^{-1}v)e^{i\sum_{j=1}^{n}(\xi_{j}-\zeta_{j})\cdot x}\mathrm{d}v.$
The first term corresponds to moments from the Itô-Schrödinger equation which was analyzed in [5, Theorem 2.7]. The result there shows that
$\sup_{s\in[0,z]}\mathbb{E}|\delta_{h}\phi^{\varepsilon}(z,r,x)|^{2n}\leq C(z,n)|h|^{2n\alpha}\,,$
where $0<\alpha\leq 1$ in the kinetic regime and $10<\alpha<$ 1 in the diffusive regime.
The second part can be bounded as follows. Performing a change of variables $\eta^{-1}v\to v$ shows that this is bounded by
$\int\limits_{\mathbb{R}^{2n}}\prod\limits_{j=1}^{n}[|e^{i\eta\xi_{j}\cdot h}-1||e^{-i\eta\zeta_{l}\cdot h}-1|]|\mathscr{R}^{\theta}_{p,q}(z,v)|\prod\limits_{j=1}^{n}\mathrm{d}\xi_{j}\mathrm{d}\zeta_{j}$ $\leq\eta^{2n}|h|^{2n}\|\prod\limits_{m=1}^{2n}\langle v_{m}\rangle\mathscr{R}^{\theta}_{n,n}(z,v)\|\,.$
Due to Theorem 4.3, we have that this is bounded by
$|h|^{2n}\theta\eta^{2n-2}e^{\frac{\lambda[R,n,n,z,1]}{\eta^{2}}}\|\prod\limits_{m=1}^{2n}\langle v_{m}\rangle^{2}\psi^{\theta}(0,v)\|,$
where from (36), $\lambda[R,n,n,z,1]=c(n,z)\|\langle k\rangle^{2+2n}\hat{R}(k)\|$. Due to assumption (9), we can find an $n$ such that $12n\geq d+$ 1 and $\|\langle k\rangle^{2+2n}\hat{R}(k)\|$ is bounded so that the term above is bounded by $|h|^{2n}$ (and thereby $|h|^{2n\alpha}$) uniformly in $\theta$.
The tightness and stochastic continuity of the process $x\to\phi^{\theta}(z,r,x)$ in $C^{0,\alpha_{-}}(\mathbb{R}^{d})$ is then a standard result; see, e.g., [24]. ∎
5 Conclusions
Starting from the paraxial approximation, we have characterized the limiting statistical distribution of high frequency wavefields propagating over long distances through weak turbulence. In the scintillation regime considered here, we show that an appropriately rescaled macroscopic wavefield is asymptotically described by a complex Gaussian distribution with scintillation saturating to unity. This is consistent with experimental observations of speckle patterns where intensity of the wavefield is modeled by an exponential law. We have used a Gaussian assumption with short range correlations on the refractive index fluctuations to simplify our derivations. This allows us to show that the statistical moments of the wavefield of the paraxial wave equation are well approximated by moments given by closed-form PDEs of the Itô-Schrödinger model. Our analysis is primarily done in the Fourier domain with errors controlled in the total variation norm which translates to error estimates in the uniform sense in the physical domain. This choice is important for applications as physical manifestations like speckle are observed predominantly in the spatial domain. This in turn allows us to use the results in [5] where the moments of the Itô-Schrödinger equation have been shown to converge to that of a complex Gaussian random variable. The extension of these results to a more general class of turbulence scenarios where the Gaussian assumption on the randomness is relaxed is left for future investigation.
Acknowledgments
We would like to thank Christophe Gomez for many stimulating discussions on the validity of the paraxial and Itô-Schrödinger models. This work was funded in part by NSF grant DMS-230641.